OpenStudy (yourboydusty):

In the square below, the four quarter-circles are congruent. Find the area of the shaded region. If necessary, round your answer to two decimal places.

OpenStudy (yourboydusty):

@3mar

OpenStudy (3mar):

What did you pick?

OpenStudy (yourboydusty):

i honestly have no clue how to do this... iwas not taught how to :(

OpenStudy (3mar):

The shaded area is the subtraction of 4 quarter-circles areas from the area of the square.

OpenStudy (3mar):

It is soo tricky and soo sooo easy in the same time

OpenStudy (3mar):

radius of circle is s area of a quarter circle is \[\frac{ 1 }{ 4 }(\pi*5^2)=\frac{ 25\pi }{ 4 }\] area of the 4 quarters will be \[\frac{ 25\pi }{ 4 }*4=25\pi\]

OpenStudy (3mar):

area of the square is 64

OpenStudy (sshayer):

area of square =x*x there are four quarters of a circle each of radius 4 area of four sectors=\[4 \pi 4^2\times \frac{ 90 }{ 360 }=16 \pi\] area of shaded region=area of square-area of 4 sectors=?

OpenStudy (yourboydusty):

....... so confused................

OpenStudy (3mar):

the shaded area =64-16pi=13.735 sorry the radius will be 4 (8/2) not 5

OpenStudy (yourboydusty):

so 4 and 8/2 is the answer?

OpenStudy (3mar):

radius of circle is s area of a quarter circle is 0.25(π∗4^2)=4π area of the 4 quarters will be 4π∗4=16π

OpenStudy (3mar):

This is the answer: the shaded area =64-16pi=13.735

OpenStudy (sshayer):

\[area ~of ~one ~sector =\pi r^2 \times \frac{ \theta }{ 180 }\] area of 4 sectors\[=4 \pi r^2\times \frac{ \theta }{ 360 }\] \[\theta=90\]

OpenStudy (3mar):

@YourBoyDusty concentrate at one of us and follow with him to the end

OpenStudy (yourboydusty):

ok....next

OpenStudy (3mar):

yes

OpenStudy (3mar):

Thank you for the medal!